# AI-driven View Guidance System in Intra-cardiac Echocardiography Imaging

### v1

@@ Abstract
Intra-cardiac Echocardiography (ICE) is a crucial imaging modality used in electrophysiology (EP) and structural heart disease (SHD) interventions, providing real-time, high-resolution views from within the heart. Despite its advantages, effective manipulation of the ICE catheter requires significant expertise, which can lead to inconsistent outcomes, particularly among less experienced operators. To address this challenge, we propose an AI-driven closed-loop view guidance system with human-in-the-loop feedback, designed to assist users in navigating ICE imaging without requiring specialized knowledge. Our method models the relative position and orientation vectors between arbitrary views and clinically defined ICE views in a spatial coordinate system, guiding users on how to manipulate the ICE catheter to transition from the current view to the desired view over time. Operating in a closed-loop configuration, the system continuously predicts and updates the necessary catheter manipulations, ensuring seamless integration into existing clinical workflows. The effectiveness of the proposed system is demonstrated through a simulation-based evaluation, achieving an 89% success rate with the 6532 test dataset, highlighting its potential to improve the accuracy and efficiency of ICE imaging procedures.
@@ 1 Introduction
The Intra-cardiac Echocardiography (ICE) is a sophisticated imaging modality that offers real-time, high-resolution views from within the heart, making it an invaluable tool in both electrophysiology (EP) and structural heart disease (SHD) interventions. The ICE is commonly used in procedures such as atrial fibrillation ablation, transcatheter valve repairs, and closures of septal defects. The key advantages of ICE include its ability to provide clear and detailed near-field images of cardiac structures, perform under conscious sedation, and be easily manipulated and interfaced with other interventional equipment. Additionally, the ICE has been shown to reduce procedural and fluoroscopy times, decrease overall radiation exposure to both the patient and physician, and shorten hospital stays. As a result, the ICE has become widely used in interventional procedures for the management of complex cardiac conditions [1, 2, 3, 4].
@@
The primary use cases of the ICE imaging involve visualizing target anatomy, detecting and tracking therapeutic devices, and validating treatments in real-time. For ICE imaging to be effective, the ICE catheter must be accurately positioned at the target location. Achieving this requires physicians to have significant expertise in interpreting anatomical views via ICE images and skillfully maneuvering the ICE catheter using two knobs (anterior-posterior, right-left) and the rotating/translating the catheter body itself to the target anatomical views, all while primarily focusing on the therapeutic device manipulation [5, 6]. This complexity can be particularly challenging for those who are less experienced. As a result, the procedure is largely dependent on the skill of the operator, which might lead to inconsistent outcomes and extended procedure times, emphasizing the importance of experience in achieving optimal results.
@@
Figure 1: Overview of the proposed view guidance system. When users want to navigate to clinically pre-defined views during the procedures, our proposed system provides continuous guidance on how to manipulate the ICE catheter via an interactive device (e.g., a touchpad for view selection and a feedback monitor) until the target view is achieved. (ACUSON Origin Cardiovascular Ultrasound System and AcuNav Lumos ICE catheter, Source: Siemens Healthineers)
@@
In this research, we propose an AI-driven closed-loop view guidance system for ICE imaging, designed to enable even novice users to effectively utilize ICE imaging without specialized knowledge. Figure 1 presents an overview of the proposed view guidance system. Specifically, our proposed method models the relative position and orientation vectors between arbitrary views and clinically defined common ICE views [7, 8, 9] — such as Right Ventricle (RV), Left Ventricle (LV), Left Atrial Appendage (LAA), Left Pulmonary Vein (LPV), Right Pulmonary Vein (RPV), or Esophagus (ESO) views — within a unified spatial coordinate system. The users are guided on how to manipulate the ICE catheter to transition from the current view to the desired view over time. The proposed model operates in a closed-loop configuration with a human in the loop: it predicts the necessary manipulations to the ICE catheter’s state between the current and goal views, and estimates the current catheter state to manipulate by comparing the differences between the current and home views, as well as the goal and home views, as the catheter is manipulated. This proposed system avoids disrupting the existing clinical workflow, making it seamlessly integrable into current practices. Finally, the system is evaluated by simulation-based demonstrator, where the boundaries of the predictions from our quantile loss were assessed with the collected clinical data.
@@ 2.1 US view navigation
To improve the clinical workflow, various research efforts have focused on developing guidance systems for tasks such as view finding and clinical decision-making. In particular, Artificial Intelligence (AI)-based methods have been widely utilized to guide and support these workflow. For instance, Narang et al. [10] proposed novel software that provides real-time prescriptive guidance to novice operators for obtaining the Trans-Thoracic Echocardiographic (TTE) images. Similarly, Sabo et al. [11] proposed a real-time guidance system to improve the echo-cardiographic acquisition process, and Pasdeloup et al. [12] developed a method to help users navigate toward preferred standard views during scanning. Additionally, Amadou et al. [13] demonstrated a synthetic image generation pipeline for navigation in an echocardiography view classification experiment. Building on this, Amadou et al. [14] proposed a framework using reinforcement learning to guide novice sonographers in navigating to standard interventional views during Transesophageal echocardiography (TEE).
While many pioneering works have significantly improved clinical workflows, the aforementioned studies primarily deal with acquiring echocardiograms externally or rely on simulation data. In contrast, our research is the first to specifically address a view guidance system for ICE catheter manipulation based on real clinical data. This approach requires a real-time understanding of internal cardiac structures, providing continuous guidance through the interactive mapping of images to spatial and catheter joint space. Furthermore, unlike simulation-based methods, which may not fully capture the complexities of real clinical environments, our method is designed for seamless integration into actual clinical workflows, effectively guiding the catheter within the heart’s intricate anatomy.
Although some robotic systems [15, 16, 17] exist to assist with catheter control, they are often expensive and complex to integrate into the clinical workflow, typically focusing on reproducing previously saved views by manual view survey [15, 16] or relying on positional sensors within a spatial coordinate system [17]. In contrast, our approach relies solely on ICE-imaging-based state estimation to guide the user in manipulating the ICE catheter within the control space from the current view to the target view, eliminating the need for positional sensors and avoiding disruption to the existing clinical workflow.
@@ 2.2 From Transformer to the Mamba and its medical applications
To address the relationship between non-linear behaviors in images and associated multi modal states, many recent studies have turned to AI, particularly leveraging a Transformer architecture [18]. The Transformer has been widely adopted in vision tasks due to its powerful ability to understand context across entire images through its self-attention mechanism, as well as its scalability to large models and robustness. As a result, many tasks that traditionally relied on CNN structures are now being addressed with Transformer-based architectures [19, 20, 21, 22, 23].
Recently, a new alternative to the transformer architecture, called Mamba [24], has emerged. Unlike the transformer architecture, which relies on self-attention mechanisms with quadratic computational complexity for long sequences, the State Space Model (SSM) maintains linear computational complexity. Mamba is based on the SSM, which can be represented as follows:
$\begin{split}h[k]&=\hat{A}h_{[}k-1]+\hat{B}x[k]\\ y[k]&=Ch[k],\end{split}$ (1)
where the $h[k]$ represents the state at the $k$-th step, and the matrix $C$ is the projection matrix that projects the current state to the output. The matrices $\hat{A}$ and $\hat{B}$ are the discretized versions of the system matrix and projection matrix, respectively, and they can be represented as follows:
@@
where the $\Delta$ denotes the timescale parameter. Mamba implements the Selective Scan Mechanism with SSM (S6), filtering out irrelevant information to effectively handle long-range sequences and induction heads. To achieve this, it allows the matrices $B$, $C$, and $\Delta$ to depend on the input sequence. Additionally, its hardware-aware implementation makes Mamba a computationally efficient model, resulting in significant performance gains and resource efficiency. Gu and Dao [24] first proposed the Mamba structure to address the computational burden of long sequences in transformer architecture. They deployed selective State Space Models (SSMs) and a hardware-aware parallel algorithm without using attention or MLP blocks.
Due to its outstanding performance, many researchers have started replacing the Transformer architecture with the Mamba structure. Recently, there have been numerous attempts to adapt the Mamba structure to the medical domain for tasks such as segmentation, classification, and detection, where it has shown superior performance compared to conventional methods. For instance, the authors in [25, 26, 27, 28, 29] proposed Mamba-based methods for medical image segmentation. Yue and Li [30] introduced a medical image classification method that combines convolutional neural networks (CNNs) with the Mamba structure. In their approach, local features are extracted using CNNs, while global features are obtained from the Mamba structure, which are then combined for classifying various medical images. In [31], the author presented MambaMorph, a Mamba-based multi-modality deformable registration framework. Additionally, Schiff et al. [32] utilized bi-directional Mamba blocks to model genomic sequences. Inspired by this pioneering work, we also applied the Mamba structure in our research. Specifically, our goal is to find and match the global features of ICE images to coordinates within a spatial coordinate system, which requires leveraging the entire content of the image. In this context, we believe that the Mamba structure offers more significant performance than Transformers, and we have integrated it into our research.
Figure 2: Overview of the proposed method. (a)Procedure Overview: The clinician first positions the ICE catheter at the home view and then selects the desired viewpoint. The proposed network provides the relative state, which is transformed into robotic coordinates using inverse kinematics (IK). Guided by this information, the clinician controls the ICE catheter, and the entire procedure can be performed iteratively. (b) Iteration step: During the procedure, the user can move sequentially toward the guided target state. To track progress from the start point to the target point, the proposed method acts as a state estimator, continuously updating the current state based on the home view state. Each step can be repeated as needed until the target point is reached.
Figure 3: (a) The constructed cardiac anatomical volume mesh (b) The reconstructed volume mesh for missing structures such as RA, RV, LV. We predicted its center of the mesh based on existing small dataset. (c) The final dataset composition in single subject. There exist several views of the images visualizing different structure in different position. (d) Our training set. We acquired the number of target states based on single image in single subject. The target state has a single position, but different orientation by aligning the fan direction to the center of the target volume.
@@ 3.1 Problem Setup and Notation
The ICE catheter is manually introduced into the femoral vein at the groin through an introducer sheath. A basic ICE study typically begins with the catheter positioned in the mid-right atrium (RA), achieving the home view with the ICE catheter in a neutral position, providing imaging of the RA, tricuspid valve, and right ventricle [33]. Clinicians often rely on this home view as a view that can be easily recovered when encountering difficulties in ICE imaging. We assume this home view as the reference coordinate system for our proposed method.
Since most ICE catheter work is conducted in the RA, our guidance system assumes that there are no obstacles in the RA, and therefore, obstacle avoidance is not addressed in this article. Additionally, we assume that ICE imaging is captured during the diastolic phase, when the heart is relaxed and the ventricles are filled with blood, particularly during atrial systole when the QRS complex is prominent [34].
Finally, while we use the tip’s position and orientation information for modeling, it is important to note that our actual guidance system does not require attaching position/orientation sensors to the ICE catheter. This allows our system to function as a purely ICE-image-based guidance system, eliminating the need for additional sensor attachments on the catheter.
We define the ICE imaging state S = ($x,y,z,\delta x,\delta y,\delta z$) as position and orientation in $\mathbb{SE}(3)$. The ICE catheter has four degrees of freedom: two DOFs for steering the catheter tip in two planes (anterior-posterior knob angle $\theta_{1}$ and the right-left knob angle $\theta_{2}$) using two knobs on the catheter handle, bulk rotation $\theta_{3}$, and translation $d_{4}$ along the major axis of the catheter. Thus, the ICE joint state, J = $(\theta_{1},\theta_{2},\theta_{3},d_{4})$. The ICE imaging, I, is defined as the acquired images, which can be either 2D or 3D.
Since the home view is a neutral position and one of the pre-defined clinical views, we have a transition function from S to J using inverse kinematics ($\mathcal{IK}$) [35] based on the home view (i.e., ${\bf S}_{home}$) as a reference coordinate system:
@@
where $i$ and $j$ represent the index of the possible ICE views.
Problem Definition: Given the current imaging ${\bf I}_{curr}$ and a goal index $g$ from a library of views, the objective is to estimate the relative ICE joint state $\Delta{\bf J}_{(curr,g)}$ over time:
@@ 3.2 Procedure Overview
Figure 2 (a) presents an overview of the proposed ICE view guidance system. The procedure is broken down into several key steps, each of which is detailed below:
@@
    Home view positioning: The procedure begins by positioning the ICE catheter in the mid-right atrium (RA) to achieve the home view. Then, the ICE catheter is placed in a neutral position, which serves as our reference coordinate system.
@@
    The desired view selection: The clinician selects the goal view (as index $g$) from a predefined library of clinical views. The selected view index is then fed into our model $\mathcal{M}$ to estimate the current imaging state in spatial space $S_{home}^{g}$ (Equation (5)).
@@
    Compute relative vectors $\Delta{\bf J}_{(curr,g)}$: This process requires the continuous prediction of the relative vector $\Delta{\bf J}_{(curr,g)}$ to provide feedback to the user. To achieve this, it is essential to know the current catheter imaging state ${\bf S}_{curr}^{home}$ (Equation (6)) to calculate the relative vectors needed for movement. A novel aspect of our approach is that our same model also serves as a state estimator for this purpose. Then, using $\mathcal{IK}$, the relative joint vectors $\Delta{\bf J}_{(curr,g)}$ from the current state to the goal state are computed at Equation (7). Figure 2 (b) illustrates the iterative process.
@@
Figure 4: The proposed method architecture. (a) The whole network architecture. The center of pink block is MedMamba structure. We added layers to feature mixing with target class code and image. At the output of the network, we added multi-head structure to estimate position and orientation separately. (b) The details of SS2D structure. The image feature is split into sequence in different order and put into SSM model. Then, the output is merged together.

### v3

@@ Abstract
Intra-cardiac echocardiography (ICE) is a crucial imaging modality used in electrophysiology (EP) and structural heart disease (SHD) interventions, providing real-time, high-resolution views from within the heart. Despite its advantages, effective manipulation of the ICE catheter requires significant expertise, which can lead to inconsistent outcomes, especially among less experienced operators. To address this challenge, we propose an AI-driven view guidance system that operates in a continuous closed-loop with human-in-the-loop feedback, designed to assist users in navigating ICE imaging without requiring specialized knowledge. Specifically, our method models the relative position and orientation vectors between arbitrary views and clinically defined ICE views in a spatial coordinate system. It guides users on how to manipulate the ICE catheter to transition from the current view to the desired view over time. By operating in a closed-loop configuration, the system continuously predicts and updates the necessary catheter manipulations, ensuring seamless integration into existing clinical workflows. The effectiveness of the proposed system is demonstrated through a simulation-based performance evaluation using real clinical data, achieving an 89% success rate with 6,532 test cases. Additionally, a semi-simulation experiment with human-in-the-loop testing validated the feasibility of continuous yet discrete guidance. These results underscore the potential of the proposed method to enhance the accuracy and efficiency of ICE imaging procedures
@@ 1 Introduction
Intra-cardiac echocardiography (ICE) is a sophisticated imaging modality that provides real-time, high-resolution views from within the heart, making it an invaluable tool in both electrophysiology (EP) and structural heart disease (SHD) interventions. ICE is commonly used in procedures such as atrial fibrillation ablation, transcatheter valve repairs, and septal defect closures. The key advantages of ICE include its ability to provide clear and detailed near-field images of cardiac structures. It can be performed under conscious sedation, manipulated easily, and interfaced with other interventional equipment. Additionally, ICE has been shown to reduce procedural and fluoroscopy times, decrease overall radiation exposure to both the patient and physician, and shorten hospital stays. As a result, ICE has become widely used in interventional procedures for the management of complex cardiac conditions [1, 2, 3, 4].
@@
The primary use cases of ICE imaging involve visualizing target anatomy, detecting and tracking therapeutic devices, and validating treatments in real-time. For ICE imaging to be effective, the ICE catheter must be accurately positioned at the target location. Achieving this requires physicians to have significant expertise in interpreting anatomical views via ICE images and skillfully maneuvering the ICE catheter using two knobs (anterior-posterior, right-left). They must also rotate and translate the catheter body itself to reach the target anatomical views, while primarily focusing on therapeutic device manipulation [5, 6]. This complexity can be particularly challenging for those who are less experienced. As a result, the procedure is largely dependent on the skill of the operator, which might lead to inconsistent outcomes and extended procedure times. It emphasizes the importance of experience in achieving optimal results.
@@
Figure 1: Overview of the proposed view guidance system. When users want to navigate to clinically predefined views during procedures, our proposed system provides continuous guidance on how to manipulate the ICE catheter via an interactive device (e.g., a touchpad for view selection and a feedback monitor) until the target view is achieved. (ACUSON Origin Cardiovascular Ultrasound System and AcuNav Lumos ICE catheter, Source: Siemens Healthineers)
@@
In this research, we propose an AI-driven view guidance system for continuous closed-loop guidance with human-in-the-loop feedback, designed to assist users in manipulating ICE imaging. The system aims to empower even novice users to effectively operate ICE imaging without requiring specialized expertise. Figure 1 presents an overview of the proposed view guidance system. Specifically, our proposed method models the relative position and orientation vectors between arbitrary views and clinically defined common ICE views [7, 8, 9] - such as the Right Ventricle (RV), Left Ventricle (LV), Left Atrial Appendage (LAA), Left Pulmonary Vein (LPV), Right Pulmonary Vein (RPV), or Esophagus (ESO) views-within a unified spatial coordinate system. The proposed model can be utilized in a closed-loop configuration with a human-in-the-loop, predicting the necessary manipulations to the ICE catheter’s state between the current view and a clinically defined target view (i.e., the goal view). The user is guided step-by-step to perform the required manipulations on the ICE catheter, with continuous instructions provided until the desired view is successfully achieved.
@@ 2.1 US view navigation
To improve clinical workflow, various research efforts have focused on developing guidance systems for tasks such as view finding and clinical decision-making. In particular, Artificial Intelligence (AI)-based methods have been widely utilized to guide and support these workflows. For instance, Narang et al. [10] proposed novel software that provides real-time prescriptive guidance to novice operators for obtaining the Trans-Thoracic Echocardiographic (TTE) images. Similarly, Sabo et al. [11] proposed a real-time guidance system to improve the echo-cardiographic acquisition process. Pasdeloup et al. [12] developed a method to help users navigate toward preferred standard views during scanning. Additionally, Amadou et al. [13] demonstrated a synthetic image generation pipeline for navigation in an echocardiography view classification experiment. Building on this, Amadou et al. [14] proposed a framework using reinforcement learning to guide novice sonographers in navigating to standard interventional views during Transesophageal echocardiography (TEE).
While many pioneering works have significantly improved clinical workflows, the aforementioned studies primarily focus on acquiring echocardiograms externally or rely on simulation data. In contrast, our research is the first to specifically address a view guidance system for the ICE catheter manipulation based on real clinical data. This approach requires real-time understanding of internal cardiac structures, providing continuous guidance through the interactive mapping of images to spatial and catheter joint spaces. Furthermore, unlike simulation-based methods, which may not fully capture the complexities of real clinical environments, our method is designed for seamless integration into actual clinical workflows, effectively guiding the catheter within the heart’s intricate anatomy.
Although some robotic systems [15, 16, 17] exist to assist with catheter control, they are often expensive and complex to integrate into the clinical workflow. These systems typically focus on reproducing previously saved views by manual view survey [15, 16] or rely on positional sensors within a spatial coordinate system [17]. In contrast, our approach relies solely on ICE-imaging-based state estimation to guide the user in manipulating the ICE catheter within the control space, from the current view to the target view. This eliminates the need for positional sensors and avoids disruption to the existing clinical workflow.
@@ 2.2 From Transformer to the Mamba and its medical applications
To address the relationship between non-linear behaviors in images and associated multi-modal states, many recent studies have turned to AI, particularly leveraging a Transformer architecture [18]. The transformer has been widely adopted in vision tasks due to its powerful ability to understand context across entire images through its self-attention mechanism, as well as its scalability to large models and robustness. As a result, many tasks that traditionally relied on convolutional neural network (CNN) structures are now being addressed with transformer-based architectures [19, 20, 21, 22, 23].
Recently, a new alternative to the transformer architecture, called Mamba [24], has emerged. Unlike the transformer, which relies on self-attention mechanisms with quadratic computational complexity for long sequences, the State Space Model (SSM) maintains linear computational complexity. The Mamba is based on the SSM, which can be represented as follows:
$\begin{split}h[k]&=\hat{A}h[k-1]+\hat{B}x[k]\\ y[k]&=Ch[k],\end{split}$ (1)
where the $h[k]$ represents the state at the $k$-th step, and the matrix $C$ is the projection matrix that projects the current state to the output. The matrices $\hat{A}$ and $\hat{B}$ are the discretized versions of the system matrix and the projection matrix, respectively. They can be represented as follows:
@@
where the $\Delta$ denotes the time-scale parameter.
Mamba implements the Selective Scan Mechanism with SSM (S6), filtering out irrelevant information to effectively handle long-range sequences and induction heads. To achieve this, it allows the matrices $B$, $C$, and $\Delta$ to depend on the input sequence. Additionally, its hardware-aware implementation makes Mamba computationally efficient model, leading to significant performance gains and improved resource efficiency.
Due to its outstanding performance, many researchers have started replacing the transformer architecture with the Mamba structure. Recently, numerous efforts have been made to adapt the Mamba structure to the medical domain for tasks such as segmentation, classification, and detection. For instance, the authors in [25, 26, 27, 28, 29] proposed Mamba-based methods for medical image segmentation. Yue and Li [30] introduced a medical image classification method that combines CNN with the Mamba structure. In their approach, local features are extracted using CNN, while global features are obtained from the Mamba structure, which are then combined to classify various medical images. In [31], the author presented MambaMorph, a Mamba-based multi-modality deformable registration framework. Additionally, Schiff et al. [32] utilized bi-directional Mamba blocks to model genomic sequences.
Our goal is to find and match the global features of ICE images to coordinates within a spatial coordinate system, which requires leveraging the entire content of the image. In this context, we believed that the Mamba structure offers superior performance compared to transformers, and we have integrated it into our research.
Building on these pioneering works, we propose a novel ICE view guidance system with the following key contributions:
  1. 1.
    We introduce a new ICE view guidance system that enables precise navigation from arbitrary views to clinically defined target views.
  2. 2.
    Our approach leverages a variation of the Mamba regression model, enabling robust and accurate guidance.
  3. 3.
    We validated the feasibility of the proposed system through comprehensive simulation and semi-simulation experiments, effectively demonstrating its real-world applicability.
@@ 3.1 Problem Setup and Notation
The ICE catheter is manually inserted into the femoral vein at the groin through an introducer sheath. A basic ICE study typically begins with the catheter positioned in the mid-right atrium (RA), achieving the home view, with the catheter in a neutral position. It provides imaging of the RA, tricuspid valve, and right ventricle [33]. Clinicians often rely on this home view as a reference that can be easily restored if difficulties arise during ICE imaging. We assume this home view as the reference coordinate system for our proposed method.
Since most ICE catheter procedures are conducted in the RA, our guidance system assumes that there are no obstacles in the RA, and therefore, obstacle avoidance is not addressed in this article. Based on the physiological properties of the cardiac cycle, we assume that ICE imaging is captured during the diastolic phase, specifically at the end of the T-wave on the ECG. This timing corresponds to when the heart is fully relaxed, the ventricles are filled with blood, and the Left Atrium (LA) reaches its maximum extent. While this assumption aligns with common settings in systems like CARTO [34] (Detailed in Section 3.3), it provides a stable and consistent reference point for imaging, minimizing motion artifacts [35].
Finally, while we use the tip’s position and orientation information for modeling, it is important to note that our actual guidance system does not require position/orientation sensors attached to the ICE catheter. This allows our system to operate as a purely ICE-image-based guidance system, eliminating the need for additional sensor attachments on the catheter.
We define the ICE imaging state S = ($x,y,z,\delta x,\delta y,\delta z$) as the position and orientation in $\mathbb{SE}(3)$. The ICE catheter has four degrees of freedom: two DOFs for steering the catheter tip in two planes (anterior-posterior knob angle $\theta_{1}$ and right-left knob angle $\theta_{2}$) using two knobs on the catheter handle, bulk rotation $\theta_{3}$, and translation $d_{4}$ along the major axis of the catheter. Thus, the ICE joint state is J = $(\theta_{1},\theta_{2},\theta_{3},d_{4})$. The ICE imaging, I, is defined as the acquired images, which can be either 2D or 3D.
Since the home view is a neutral position and one of the predefined clinical views, we define a transition function from S to J using inverse kinematics ($\mathcal{IK}$) [36], with the home view (i.e., ${\bf S}_{home}$) serving as the reference coordinate system:
@@
where $j$ represents the index of the possible ICE views.
Problem Definition: Given the current imaging ${\bf I}{curr}$ and a goal index $g$ from a library of views, the objective is to estimate the relative ICE joint state, $\Delta{\bf J}{(curr,g)}$, over time:
@@
Figure 2: Process of the proposed method. (a)Procedure Overview: The clinician first positions the ICE catheter at the home view and then selects the desired viewpoint. The proposed network provides the relative state, which is then transformed into robotic coordinates using inverse kinematics (IK). Guided by this information, the clinician manipulates the ICE catheter, and the procedure is performed iteratively. (b) Iteration step: During the procedure, the user can sequentially move toward the guided target state. The proposed method continuously updates the current state based on the home view state, enabling progress tracking from the starting point to the target point. Each step can be repeated as needed until the target view is achieved.
Figure 3: (a) The reconstructed cardiac anatomical volume mesh of a single subject, generated using our volume contouring algorithm [37]. (b) The reconstructed volume mesh for missing structures such as the RA, RV, and LV, where the mesh centers were predicted based on a small existing dataset. (c) The composition of the CARTO dataset for a single subject, which includes multiple image views visualizing different structures in various positions, along with corresponding position/orientation data and the volume mesh. (d) Our target dataset consists of six target states, each corresponding to one of six clinically defined views for a single subject. While each target state shares the same position, the orientations differ, with the fan direction aligned to the center of the target volume.
@@ 3.2 Procedure Overview
Figure 2 (a) provides an overview of the proposed ICE view guidance system. The procedure is divided into several key steps, each of which is detailed below:
@@
    Home view positioning: The procedure begins by positioning the ICE catheter in the mid-RA to achieve the home view. Next, the ICE catheter is placed in a neutral position, which serves as the reference coordinate system.
@@
    The desired view selection: The clinician selects the goal view (indexed as $g$) from a predefined library of clinical views. The selected view index is then fed into our model $\mathcal{M}$ to estimate the current imaging state in spatial space $S_{home}^{g}$ (Equation (5)).
@@
    Compute relative vectors $\Delta{\bf J}_{(curr,g)}$: This process requires the continuous prediction of the relative vector $\Delta{\bf J}_{(curr,g)}$ to provide feedback to the user. To achieve this, it is essential to know the current catheter imaging state ${\bf S}_{curr}^{home}$ (Equation (6)) to calculate the relative vectors needed for movement. Then, using $\mathcal{IK}$, the relative joint vectors $\Delta{\bf J}_{(curr,g)}$ from the current state to the goal state are computed at Equation (7). Figure 2 (b) shows the iterative process.
@@
Figure 4: The proposed method architecture. The center of the block represents the MedMamba structure. We have added layers for feature mixing with target class code and image. At the output of the network, we added multi-head structure to estimate position and orientation separately. The details of SS2D structure are shown at the bottom. The image feature is split into sequences in different orders and put into the SSM model. Then, the outputs are merged together.